# EditRoom: LLM-parameterized Graph Diffusion for Composable 3D Room Layout Editing

### v1

@@ Abstract
Given the steep learning curve of professional 3D software and the time-consuming process of managing large 3D assets, language-guided 3D scene editing has significant potential in fields such as virtual reality, augmented reality, and gaming. However, recent approaches to language-guided 3D scene editing either require manual interventions or focus only on appearance modifications without supporting comprehensive scene layout changes. In response, we propose EditRoom, a unified framework capable of executing a variety of layout edits through natural language commands, without requiring manual intervention. Specifically, EditRoom leverages Large Language Models (LLMs) for command planning and generates target scenes using a diffusion-based method, enabling six types of edits: rotate, translate, scale, replace, add, and remove. To address the lack of data for language-guided 3D scene editing, we have developed an automatic pipeline to augment existing 3D scene synthesis datasets and introduced EditRoom-DB, a large-scale dataset with 83k editing pairs, for training and evaluation. Our experiments demonstrate that our approach consistently outperforms other baselines across all metrics, indicating higher accuracy and coherence in language-guided scene layout editing.
@@ 1 Introduction
Therefore, we propose EditRoom, a unified framework that can execute all editing types by complex natural language commands without intermediate manual interventions. It consists of two main modules: the command parameterizer and the scene editor. The command parameterizer employs an pretrained LLM, specifically GPT-4o (OpenAI, 2024), to transform natural language commands into sequences of breakdown commands for six basic editing types on single object: adding, removing, replacing, translating, rotating, and scaling. These breakdown commands, along with the source scenes, are then fed sequentially into the scene editor for execution. To unify all editing types, we convert scenes into graph representations and construct scene editor as conditional graph generation models, where we take source scenes and text commands as conditions and train diffusion models to generate the target scene graphs with the layout.
@@ 2 Related Work
Language-guided 3D Scene Generation  Currently, some language-guided 3D Scene generation works that claim for language-guided scene editing capablility can be abstractly categorized into two main approaches. The first approach starts with generating the scene configurations by utilizing large language models, and prompt large language model for following editing (Vilesov et al., 2023; Zhou et al., 2024b; Aguina-Kang et al., 2024). However, they are limited to editing its own generated scenes and do not accept source scenes as input. The second approach is about training diffusion-based model to learn text-conditional scene generation (Haque et al., 2023; Tang et al., 2023; Zhai et al., 2024). During the editing, they needs to manually mask the editing part and leverage the completion ability of diffusion models. Besides, their editing types are limited. InstructScene (Haque et al., 2023) does not add language conditions to the scene layout, so translating, rotating, and scaling cannot be applied to the scenes. EchoScene’s manipulations only apply to object relations so that they can only add an object or change the relative positions of objects (Zhai et al., 2024). In contrast to these works, EditRoom can accept an existing scene as input and apply free-form editing commands for 3D scene layout without manual interventions.
Language-guided 3D Scene Editing  Recent language-guided 3D scene editing works mostly incoprate neural field representations and pretrained image generation models for object appearance editing. Some works are mainly focusing on replacing single object (Zhuang et al., 2023; Bartrum et al., 2024; Karim et al., 2023) but failing to edit the layout. Other approaches (Chen et al., 2023; Ye et al., 2023) leverage pretrained segmantation models to obtain the individual object representation so that they can remove the objects by manually selecting the target objects. Besides, all these methods can only take one single-operation command at each interaction. Compared to these previous works, EditRoom can leverage LLM to deal with multi-operation natural language commands and automatically execute all editing types through a unified graph diffusion-based module.
@@ 4 The EditRoom-DB Dataset
Our data generation pipeline supports a variety of editing operations, including Add and Remove Objects, Pose and Size Changes, and Object Replacement. Each modification produces a new scene paired with a detailed textual description of the changes made, following a predefined template. For each scene, objects are selected randomly and modified iteratively through basic editing operations. In the case of Add and Remove Objects, the modified scene serves as the target for the operation, with the original scene acting as the source. For Pose and Size Changes, random adjustments are applied to selected objects, and collision checking ensures that the final scenes are free from object overlap. Similarly, in Object Replacement, new objects from the same category are substituted for existing ones, with collision checks ensuring high-quality outputs. More details about automatic pipelines can be found in Appendix B.
@@
The resulting dataset consists of approximately 83,000 training samples and 1,500 test samples (randomly sampled 500 for each room type). Table 1 provides a detailed breakdown of the dataset statistics across the three scene categories. Each sample includes a source scene, an edited target scene, and corresponding language commands. This comprehensive dataset enables the development and evaluation of robust models capable of performing various scene editing tasks. By simulating realistic workflows and providing detailed text commands, EditRoom-DB serves as a valuable resource for advancing language-guided 3D scene editing.
@@
Translate 8.6k 3.2k 2.7k 61 58 74
Rotate 4.0k 1.3k 1.3k 38 35 27
Scale 12.7k 4.5k 3.9k 146 144 162
Add 8.9k 3.4k 2.8k 75 79 57
Remove 8.8k 3.3k 2.8k 129 142 127
Replace 6.8k 2.2k 2.1k 51 42 53
Total 49.8k 17.9k 15.6k 500 500 500
@@ 5.1 Experimental Setup
Baselines  To the best of the authors’ knowledge, EditRoom is the first work that can automatically execute all editing types with natural language commands. Therefore, we construct two baselines by modifying language-guided 3D scene synthesis methods for comparisons: DiffuScene-E and SceneEditor-N:
@@ 5.2 Results
DiffuScene-N $0.6213$ $0.6122$ $0.1374$ $0.9550$ $0.4484$ $0.4338$ $0.1984$ $0.9247$ $0.4693$ $0.4507$ $0.1748$ $0.9328$
SceneEditor-N $0.7254$ $0.7150$ $0.1071$ $0.9601$ $0.5189$ $0.5033$ $0.1572$ $0.9356$ $0.4797$ $0.4667$ $0.1638$ $0.9385$
EditRoom 0.7435 0.7344 0.0967 0.9644 0.5246 0.5095 0.1489 0.9450 0.4801 0.4724 0.1564 0.9463
@@
DiffuScene-N $0.5237$ $0.5115$ $0.1691$ $0.9488$ $0.5902$ $0.5770$ $0.1372$ $0.9510$
SceneEditor-N $0.5611$ $0.5491$ $0.1488$ $0.9511$ $0.6269$ $0.6146$ $0.1313$ $0.9526$
EditRoom 0.5782 0.5673 0.1432 0.9553 0.6277 0.6158 0.1290 0.9538
@@
DiffuScene-N $0.5816$ $0.5691$ $0.1248$ $0.9510$ $0.5662$ $0.5435$ $0.1439$ $0.9336$
SceneEditor-N $0.6191$ $0.6083$ $0.1150$ $0.9573$ $0.6002$ $0.5746$ $0.1398$ $0.9390$
EditRoom 0.6309 0.6216 0.1083 0.9610 0.6114 0.5837 0.1369 0.9427
@@
DiffuScene-N $0.5075$ $0.4946$ $0.1826$ $0.9381$ $0.4291$ $0.4144$ $0.2060$ $0.9325$
SceneEditor-N $0.5595$ $0.5472$ $0.1619$ $0.9341$ $0.5545$ $0.5412$ $0.1442$ $0.9496$
EditRoom 0.5657 0.5542 0.1571 0.9430 0.5556 0.5453 0.1363 0.9517
@@
Single-operation Commands  To assess model performance on single operations, we test our model and baselines using the EditRoom-DB test set, which contains 500 samples per room type, with language commands generated by GPT-4o. Quantitative results are depicted in Tables 2 and 3, and qualitative outcomes are illustrated in Figure 3. Table 2 indicates that EditRoom consistently outperforms other baselines across all room types, with notably superior performance in bedrooms. According to Table 3, EditRoom also excels across all editing types. Comparisons between EditRoom and SceneEditor-N reveal that template-based instructions can simplify the learning process by more effectively aligning language commands with scene changes. Moreover, the LLM (GPT-4o) demonstrates a successful bridge between natural language and template commands. SceneEditor-N outperforms DiffScene-E across all metrics and editing types, suggesting that our graph-based diffusion method yields more coherent and accurate editing results compared to the UNet-based approach. Thus, EditRoom provides more precise and coherent atomic editing operations from natural language commands than its counterparts.
@@
EditRoom (Concat-Text) 0.5992 0.5835 0.1325 0.9547
EditRoom (Original) 0.7435 0.7344 0.0967 0.9644
@@
EditRoom (OpenCLIP-ViT-bigG-14) 0.6970 0.6788 0.1271 0.9490
EditRoom (Original) 0.7435 0.7344 0.0967 0.9644
@@ Appendix A LLM as Command Planner
Appendix B EditRoom-DB pipeline details

### v2

@@ Abstract
Given the steep learning curve of professional 3D software and the time-consuming process of managing large 3D assets, language-guided 3D scene editing has significant potential in fields such as virtual reality, augmented reality, and gaming. However, recent approaches to language-guided 3D scene editing either require manual interventions or focus only on appearance modifications without supporting comprehensive scene layout changes. In response, we propose EditRoom, a unified framework capable of executing a variety of layout edits through natural language commands, without requiring manual intervention. Specifically, EditRoom leverages Large Language Models (LLMs) for command planning and generates target scenes using a diffusion-based method, enabling six types of edits: rotate, translate, scale, replace, add, and remove. To address the lack of data for language-guided 3D scene editing, we have developed an automatic pipeline to augment existing 3D scene synthesis datasets and introduced EditRoom-DB, a large-scale dataset with 83k editing pairs, for training and evaluation. Our experiments demonstrate that our approach consistently outperforms other baselines across all metrics, indicating higher accuracy and coherence in language-guided scene layout editing. Project website: https://eric-ai-lab.github.io/edit-room.github.io/
@@ 1 Introduction
Therefore, we propose EditRoom, a unified framework that can execute all editing types by complex natural language commands without intermediate manual interventions. It consists of two main modules: the command parameterizer and the scene editor. The command parameterizer employs an pretrained LLM, specifically GPT-4o (OpenAI, 2024), to transform natural language commands into sequences of breakdown commands for six basic editing types on single object: adding, removing, replacing, translating, rotating, and scaling. These breakdown commands, along with the source scenes, are then fed sequentially into the scene editor for execution. To unify all editing types, we convert scenes into graph representations and construct scene editor as conditional graph generation models, where we take source scenes and text commands as conditions and train diffusion models to generate the target scene graphs with the layout. By conditioning on the user’s natural language prompts, our model generates reasonable editing results that serve as a foundation for further refinement. Importantly, the scenes remain fully editable, allowing users to provide additional prompts for iterative improvements to achieve their desired outcome.
@@ 2 Related Work
Language-Guided 3D Scene Generation  Language-guided 3D scene generation works claiming editing capabilities can be categorized into two main approaches. The first approach utilizes large language models to generate scene configurations and perform edits (Vilesov et al., 2023; Zhou et al., 2024b; Aguina-Kang et al., 2024). However, these methods are limited to editing scenes they generate and cannot accept external source scenes. The second approach trains diffusion-based models for text-conditional scene generation (Haque et al., 2023; Tang et al., 2023; Zhai et al., 2024). These methods often require manual masking for edits and have limited editing functionalities. For example, InstructScene (Haque et al., 2023) lacks support for operations like translation, rotation, and scaling, while EchoScene (Zhai et al., 2024), which directly take target scene graphs as input, only modifies object relations manually and does not process textual instructions. In contrast, EditRoom accepts existing scenes and supports free-form editing commands for 3D layouts. By leveraging diffusion models trained on clean data, EditRoom enables comprehensive edits, including translating, rotating, scaling, adding, and removing objects, without manual intervention.
Language-guided 3D Scene Editing  Previously, some works (Ma et al., 2018) have explored the rule-based methods for language-guided 3D scene editing. Compared to these works, our diffusion-based method can generate diverse editing results, generalize across editing types. Recent language-guided 3D scene editing works mostly incoprate neural field representations and pretrained image generation models for object appearance editing. Some works are mainly focusing on replacing single object (Zhuang et al., 2023; Bartrum et al., 2024; Karim et al., 2023) but failing to edit the layout. Other approaches (Chen et al., 2023; Ye et al., 2023) leverage pretrained segmantation models to obtain the individual object representation so that they can remove the objects by manually selecting the target objects. Besides, all these methods can only take one single-operation command at each interaction. Compared to these previous works, EditRoom can leverage LLM to deal with multi-operation natural language commands and automatically execute all editing types through a unified graph diffusion-based module.
@@ 4 The EditRoom-DB Dataset
Our data generation pipeline supports a variety of editing operations, including Add and Remove Objects, Pose and Size Changes, and Object Replacement. Each modification produces a new scene paired with a detailed textual description of the changes made, following a predefined template. For each scene, objects are selected randomly and modified iteratively through basic editing operations. In the case of Add and Remove Objects, the modified scene serves as the target for the operation, with the original scene acting as the source. For Pose and Size Changes, random adjustments are applied to selected objects, and collision checking ensures that the final scenes are free from object overlap. Similarly, in Object Replacement, new objects from the same category are substituted for existing ones, with collision checks ensuring high-quality outputs. More details about automatic pipelines can be found in Appendix C.
@@
The resulting dataset consists of approximately 83,000 training samples and 7,800 test samples (randomly sampled across editing types for each room type). Table 1 provides a detailed breakdown of the dataset statistics across the three scene categories. Each sample includes a source scene, an edited target scene, and corresponding language commands. This comprehensive dataset enables the development and evaluation of robust models capable of performing various scene editing tasks. By simulating realistic workflows and providing detailed text commands, EditRoom-DB serves as a valuable resource for advancing language-guided 3D scene editing.
@@
Translate 8.6k 3.2k 2.7k 808 253 250
Rotate 4.0k 1.3k 1.3k 804 250 252
Scale 12.7k 4.5k 3.9k 805 251 252
Add 8.9k 3.4k 2.8k 747 232 240
Remove 8.8k 3.3k 2.8k 816 260 253
Replace 6.8k 2.2k 2.1k 820 254 253
Total 49.8k 17.9k 15.6k 4800 1500 1500
@@ 5.1 Experimental Setup
Baselines  To the best of the authors’ knowledge, EditRoom is the first work that can automatically execute all editing types with natural language commands. Therefore, we construct two baselines by modifying language-guided 3D scene synthesis methods for comparisons: DiffuScene-N and SceneEditor-N:
@@ 5.2 Results
DiffuScene-N $0.6220$ $0.6125$ $0.1431$ $0.9415$ $0.4473$ $0.4321$ $0.1921$ $0.9254$ $0.4332$ $0.4161$ $0.1810$ $0.9236$
SceneEditor-N $0.7055$ $0.6942$ $0.1261$ $0.9510$ $0.5262$ $0.5091$ $0.1557$ $0.9366$ $0.4618$ $0.4498$ $0.1738$ $0.9329$
EditRoom 0.7342 0.7236 0.1090 0.9597 0.5360 0.5196 0.1460 0.9460 0.4710 0.4629 0.1616 0.9446
@@
DiffuScene-N $0.5363$ $0.5279$ $0.1715$ $0.9484$ $0.6010$ $0.5913$ $0.1307$ $0.9511$
SceneEditor-N $0.5983$ $0.5905$ $0.1534$ $0.9504$ $0.6655$ $0.6556$ $0.1202$ $0.9560$
EditRoom 0.6148 0.6078 0.1388 0.9584 0.6764 0.6673 0.1067 0.9610
@@
DiffuScene-N $0.6057$ $0.5947$ $0.1250$ $0.9566$ $0.5994$ $0.5754$ $0.1488$ $0.9305$
SceneEditor-N $0.6754$ $0.6650$ $0.1105$ $0.9603$ $0.6451$ $0.6193$ $0.1418$ $0.9389$
EditRoom 0.6881 0.6788 0.1008 0.9647 0.6549 0.6305 0.1384 0.9419
@@
DiffuScene-N $0.5312$ $0.5198$ $0.1799$ $0.9263$ $0.4626$ $0.4511$ $0.1877$ $0.9281$
SceneEditor-N $0.5863$ $0.5745$ $0.1620$ $0.9355$ $0.6160$ $0.6065$ $0.1290$ $0.9489$
EditRoom 0.6001 0.5887 0.1588 0.9410 0.6399 0.6316 0.1150 0.9583
@@
Single-operation Commands  To assess model performance on single operations, we test our model and baselines using the EditRoom-DB test set, with language commands generated by GPT-4o. Quantitative results are depicted in Tables 2 and 3, and qualitative outcomes are illustrated in Figure 3. Table 2 indicates that EditRoom consistently outperforms other baselines across all room types, with notably superior performance in bedrooms. According to Table 3, EditRoom also excels across all editing types. Comparisons between EditRoom and SceneEditor-N reveal that template-based instructions can simplify the learning process by more effectively aligning language commands with scene changes. Moreover, the LLM (GPT-4o) demonstrates a successful bridge between natural language and template commands. SceneEditor-N outperforms DiffScene-E across all metrics and editing types, suggesting that our graph-based diffusion method yields more coherent and accurate editing results compared to the UNet-based approach. Thus, EditRoom provides more precise and coherent atomic editing operations from natural language commands than its counterparts.
@@
EditRoom (Concat-Text) 0.5855 0.5746 0.1432 0.9488
EditRoom (Original) 0.7342 0.7236 0.1090 0.9597
@@
EditRoom (OpenCLIP-ViT-bigG-14) 0.6882 0.6743 0.1305 0.9488
EditRoom (Original) 0.7342 0.7236 0.1090 0.9597
@@ Appendix A LLM as Command Planner
Appendix B Diffusion Denoising Process
In order for a better illustration of target scene layout generation process, we visualize intermediate steps during the diffusion denoising process for an adding operation example, shown in Fig. 5. At the start of the layout diffusion process, the poses and scales of all objects inside the target scene begin as random noise. As the diffusion process progresses, each object’s properties are iteratively refined and placed into an appropriate configuration that aligns with the source scene and the given commands. This iterative refinement highlights the strength of the diffusion model in handling scene edits cohesively. By adopting a unified approach, the diffusion model is capable of jointly generating the entire scene layout for all editing types, ensuring consistency and coherence across diverse operations.
Figure 5: Visualization of layout diffusion denoising process. The whole diffusion process is conditioned on both source scene and language commands. At the beginning of the process, the target scene layout starts from random noises. After the iterative diffusion denoising process, the target scene layout becomes coherent to source scene and command.
Appendix C EditRoom-DB pipeline details